# KNN-MMD: Cross Domain Wireless Sensing via Local Distribution Alignment

### v3

@@ III-C Model Structure
After batch normalization, the processed data is fed into ResNet to obtain the embedding, which serves as the feature space mapping mentioned in Section III-B. Finally, the MLP acts as a classifier, taking the embedding as input and producing the corresponding label as output. The network can be expressed as Eq. 10.
@@ III-D1 Preliminary Classification
    $[X^{support},Y^{support}]$: Support Set
    $[X^{test}]$: Testing Set
@@
[…]
@@
3:     $N\leftarrow\text{KNN}(x_{i},X^{support},k)$triangleright Find $k$ nearest neighbors in support set
4:     $\hat{y}_{i}\leftarrow\text{majorityVote}(Y^{support}_{N})$triangleright Predict pseudo-label
@@
8:     $c_{i}\leftarrow\frac{1}{d_{i}+\epsilon}$triangleright Compute confidence score
[…]
@@
[…]
12:$n\leftarrow\lfloor p\%\times|X^{test}|\rfloor$triangleright Number of samples to select
@@
III-D2 Training Network with Local Alignment
In Step 2, we train our network using the loss function consisting of three components, as presented in Eq. 11:
@@ III-D2 Training Network with Local Alignment
where $\alpha_{1},\alpha_{2}$ represent the weights. In the loss function, firstly, $L_{cls}$ represents the traditional cross-entropy loss function, which utilizes data solely from the training set. It is used to train the model to learn the basic classification capacity. Secondly, $L_{MMD}^{local}$ represents the local MMD, which calculates the MK-MMD between the help set and the training set within each category, as shown in Eq. (12):
@@
where $M$ represents the total category amounts, $E$ is the embedding result of ResNet, $K$ is the given kernel list, and $Y$ is the sample label. It’s used to realize our local alignment. Additionally, we add an extra global MMD, which computes the MK-MMD between the entire help set and training set, following the conventional MMD approach, as represented in Eq. (13):
@@
The early stop strategy is outlined in Algorithm 2. During the training, we keep track of the highest accuracy and the lowest value of the loss function in the support set. Each time the accuracy or the loss value is improved, we save the corresponding model parameters. When the epoch reaches the minimum training epoch $e^{min}$, if the validation accuracy does not improve and the validation loss does not decrease for consecutive $e^{threshold}$ epochs, the model is regarded as converged. Specifically, in our method, when we reach the epoch $e^{min}$, we incorporate two different factors to decrease the historical highest accuracy and increase the historical largest loss. This is because in our experiments, the model often achieves an artificially high accuracy and low loss at the beginning of training, which is not a desirable outcome. Therefore, by utilizing these factors, we aim to increase the likelihood of the neural network preserving the best parameters after reaching $e^{min}$.
@@
    $[X^{train},Y^{train}]$: Training data and labels
    $[X^{support},Y^{support}]$: Support set for domain adaptation
    $[X^{help},Y^{help}]$: Help set for validation
@@
3:$loss^{best}\leftarrow\infty$, $acc^{best}\leftarrow 0$, $\triangleright$ Initialize best validation loss and best validation accuracy
4:$e^{acc},e^{loss}\leftarrow 1$triangleright Set Epoch counters
5:$\theta^{best}\leftarrow\theta$triangleright Store best parameters
@@
8:     $E\leftarrow\text{ResNet}(X^{train})$triangleright Extract features
9:     $\hat{Y}\leftarrow\text{MLP}(E)$triangleright Compute predictions
@@
11:     $l^{train}\leftarrow\mathcal{L}(\hat{Y},Y^{train})$triangleright Eq. 11
12:     $\theta\leftarrow\theta-\eta\nabla_{\theta}l^{train}$triangleright Parameter update
13:     Validation on Help Set:
14:     $l\leftarrow\text{CrossEntropy}(X^{help},Y^{help})$triangleright Validation loss
15:     $a\leftarrow\text{Accuracy}(X^{help},Y^{help})$triangleright Validation accuracy
@@
20:         $\theta^{best}\leftarrow\theta$triangleright Save improved parameters
@@
28:         $\theta^{best}\leftarrow\theta$triangleright Save improved parameters
@@
34:         $loss^{best}\leftarrow loss^{best}\cdot\alpha$triangleright Relax loss threshold
35:         $acc^{best}\leftarrow acc^{best}\cdot\beta$triangleright Relax accuracy threshold
36:         $e^{acc},e^{loss}\leftarrow 1$triangleright Reset counters
@@
43:Return $\theta^{best}$triangleright Return best parameters found
IV Experiment
Figure 7: Hardware: Our system utilizes a commercial Wi-Fi router as the transmitter (TX) and an ESP32-S3 as the receiver (RX). The ESP32-S3 decodes the received Channel State Information (CSI) and transmits it to a PC for inference.
IV-A Dataset Description
In our experiment, we utilize two datasets: the publicly available WiGesture dataset [3] and a novel dataset named WiFall, which is proposed along with this paper. Both datasets are collected using the ESP32-S3 device and share the same data format, with a sample rate of 100Hz, 1 antenna, and 52 subcarriers. For our experiment, we divide the data into 1-second samples, resulting in each sample having a length of 100 data points.
After training, the model can be integrated into a real-time system, as shown in Fig. 7. In this system, a commercial Wi-Fi router continuously sends ping reply packets to the ESP32-S3. The ESP32-S3 extracts the CSI and forwards it to a PC equipped with an NVIDIA RTX 3090 GPU. The PC can then perform real-time inference using the trained neural network.
IV-A1 WiGesture Dataset
@@ IV-B Experiment Setup
Selected Proportion in Support Set : $p\%$ 50%
@@
$[e^{min},e^{max},e^{threshold},\alpha,\beta]$
@@
To evaluate the efficiency of our method, we compare it with several representative cross-domain methods in the fields of Wi-Fi sensing and machine learning. For fairness, we have made every effort to replicate other works according to the descriptions provided in their respective papers. For any missing settings in these papers, we adopt the same settings as our method.
Wi-Fi sensing methods can be classified into two types: signal processing methods (which extract domain-independent information) and feature learning methods [43]. Domain-independent methods typically require multiple RX-TX pairs or antennas for feature extraction (for example, WiDar3.0 [44] uses 1 TX and 6 RXs), which is not applicable in our experimental setting, where only single RX-TX pairs and a single antenna are available. For feature learning methods, most works follow a similar paradigm: they first train a feature extractor and then compare the similarity of the extracted features from the source and target domains for classification. We have selected four representative methods for comparison: Ding et al. [45], AutoFi [46], Yang et al. [47], and CrossFi [18]. Ding et al. [45] propose using a CNN trained in the source domain as a feature extractor and then use cosine similarity in the feature embedding for similarity calculations. Additionally, they employ a low-pass Butterworth filter for denoising. Based on this, AutoFi [46] is introduced (without the low-pass denoising), proposing a pre-training method to enhance the feature extractor’s capacity in the neural network. In Yang et al. [47], a different similarity calculation method is employed, specifically a Siamese network [23]. They also choose a CNN combined with Bi-LSTM as the feature extractor. Building on this, CrossFi [18] uses a ResNet-based Siamese architecture, introducing a template generation method to enhance feature representation and an attention-based similarity calculation method to better capture relationships between samples. In CrossFi, the application of the Siamese network is extended to a zero-shot scenario. In feature learning methods, the two most important factors are the feature extractor and the similarity calculation method. We selected these four methods because they represent different powerful feature extractors and similarity calculation techniques.
In the field of Wi-Fi sensing, most works are based on the feature learning paradigm mentioned above. Even some works combined with Domain Adaptation Learning (DAL) methods, such as Yang et al. [47] and CrossFi [18], which also utilize the MK-MMD loss, do not rely solely on it. To illustrate the performance of the DAL methods, we have chosen several popular methods in machine learning, including adversarial learning-based DANN [28] and ADDA [48], as well as feature-space alignment methods such as MK-MMD [40] and GFK [26]. Since these methods do not depend on specific network structures, we use the same ResNet architecture as in this paper. Additionally, we employ the same kernel functions in MK-MMD as in our KNN-MMD method. Furthermore, we also compare the methods of Tian et al. [36] and EEG [37], which also discuss local alignment methods, with detailed descriptions provided in Section II-C. All these DAL methods fall under the zero-shot category, meaning they only require a labeled training set from the source domain and an unlabeled testing set, without the necessity for a labeled support set.
IV-C Experiment Result
@@
IV-D Comparison with Traditional DA Methods
@@
IV-D1 Metric-based Method
For metric-based methods like KNN [21], the biggest problem is that they are heavily influenced by the quality of the support set. As shown in Table VI, we reduced the data dimension in the target domain to $d$ using UMAP and tested the KNN performance in an $n$-shot scenario with $k$ nearest neighbors. Each scenario was tested 3 times, and the table shows the accuracy range. It can be seen that the accuracy fluctuates greatly, which may pose challenges for the practical application of this method. In Fig. 10, it can be seen that our KNN-MMD method does not exhibit such large fluctuations. Moreover, in Table VI, it’s noticeable that our KNN-MMD is not very sensitive to the hyperparameters.
@@ IV-D1 Metric-based Method
IV-D2 Learning-based Method
@@
IV-D3 DAL Method
As mentioned before, the global alignment operation of traditional DAL methods cannot guarantee the model performance. In Fig. 12, we illustrate the UMAP result of different methods, including:
  • $\bullet$
    ResNet (no alignment): We directly train the ResNet18 on the source domain and use it to embed samples from the target domain.
  • $\bullet$
    MK-MMD (global alignment): Building on the ResNet (no alignment), we incorporate the external MK-MMD loss function to achieve global alignment between the source and target domains.
  • $\bullet$
    KNN-MMD (shot number=1): This method, proposed in this paper, aligns the source and target domains locally (within each category) by utilizing source domain data with real labels and help set data with pseudo labels.
  • $\bullet$
    Theoretical Upper Bound (local alignment): This method, based on KNN-MMD, uses the ground truth labels from the target domain to achieve local alignment. It represents the upper bound for our KNN-MMD approach, which can be realized when all pseudo labels are accurate.
It can be seen that the performance of the global alignment method MK-MMD does not differ much from ResNet, which even has no alignment operation. Even though they can split samples in the source domain (training set) well, achieving an accuracy over 99%, they cluster samples in the target domain (testing set) chaotically. In contrast, our KNN-MMD still performs well in the target domain. As shown in Fig. 12c, it successfully aligns samples from the source domain and target domain within each category. Furthermore, in Fig. 12d, we show the upper bound performance of KNN-MMD, where we provide the ground truth label of the help set instead of pseudo-labels. We notice that the cluster performance and accuracy of KNN-MMD and the upper bound do not have a significant difference, which proves the efficiency of our help set construction method. Finally, even though our early stop method can be used in most DAL methods, as the support set does not directly participate in training, limited by the model principle, even if we make it stop at the point with the best testing accuracy, it still has a poor performance, as shown in Fig. 9.
@@ IV-E Ablation Study
IV-F Expanded Experiment
IV-F1 Evaluation on WiDar3.0
@@ IV-F Expanded Experiment
To further illustrate the efficiency of our KNN-MMD, we conduct cross-domain experiments using the WiDar3.0 dataset [44], which includes both cross-people and cross-room gesture recognition scenarios. In WiDar3.0, the Body-coordinate Velocity Profile (BVP) extracted from CSI is provided, with the dimension of $t\times 20\times 20$, where $t$ represents the time length. To standardize the data dimensions, we utilize only the first 28 frames of each sample and filter out those samples that are shorter than 28 frames. For our experiment, we select the subsets as follows:
  • $\bullet$
    Cross-people Scenario: For the cross-person gesture task, we utilize data collected on November 9, 2018, which includes six gestures: Push&Pull, sweep, clap, slide, draw, zigzag, and draw-N. We designate User 1 as the source domain and User 2 as the target domain. Both the source and target domains contain 3,000 samples.
  • $\bullet$
    Cross-room Scenario: For the cross-room gesture task, we utilize data collected on November 21 and November 27, 2018, from User 2, which includes six gestures: slide, draw-O, draw zigzag, draw-N, draw-triangle, and draw-rectangle. We designate Room 1 as the source domain and Room 2 as the target domain. The source domain contains 3,000 samples, while the target domain has 750 samples.
Figure 13: Experiment Results on WiDar3.0 [44]: The solid line represents the performance of KNN-MMD, while the dashed and dotted lines indicate the benchmark results for cross-room and cross-person scenarios, respectively.
The experimental results are shown in Fig. 13, where our KNN-MMD demonstrates excellent performance in both tasks, significantly outperforming the benchmark. This benchmark consists of ResNet18 [20], which is directly trained using the BVP features from the source domain without utilizing any data from the target domain.
IV-F2 Influence of Model Input
In many Wi-Fi sensing studies, processed CSI signals, such as Time of Flight (ToF), Angle of Arrival (AoA), Attenuation, and Dynamic Frequency Selection (DFS), are commonly used instead of the original CSI. These processed signals serve as important feature extraction methods in model-based approaches. However, this practice remains controversial in the field of deep learning. While data pre-processing methods can help extract meaningful features, they may also lead to the loss of information present in the original data. As noted in [44], DFS contains the most information about velocity distribution, offering advantages over other features. To investigate this, we compare the performance of KNN-MMD on the WiGesture dataset using both original CSI signals and DFS, as illustrated in Fig. 14. The results indicate that DFS can improve model performance in the task of person identification, but may hinder action recognition. A potential reason for this is that DFS is correlated with the location and orientation of individuals. In the WiGesture dataset, different people perform actions in varying directions, which creates a significant distribution gap between the source and target domains, even when using DFS. Conversely, for the same person, different actions are conducted in similar directions and at similar speeds, allowing DFS to be beneficial. (During data collection, volunteers performed different gestures continuously, and data was collected in turn.) However, we observe that the performance gap between the two types of input is relatively small, approximately 5%, which is acceptable for practical applications.

### v4

@@ IV-C Model Structure
After batch normalization, the processed data is fed into ResNet to obtain the embedding, which serves as the feature space mapping mentioned in Section IV-B. Finally, the MLP acts as a classifier, taking the embedding as input and producing the corresponding label as output. The network can be expressed as:
@@ IV-D1 Preliminary Classification
$X^{support},Y^{support}]$: Support Set
$X^{test}]$: Testing Set
@@
1:Initialize empty list $\mathit{confidences}\leftarrow\emptyset$
@@
3:  $N\leftarrow\text{KNN}(x_{i},X^{support},k)$ $\triangleright$ Find $k$ nearest neighbors in support set
4:  $\hat{y}_{i}\leftarrow\text{majorityVote}(Y^{support}_{N})$ $\triangleright$ Predict pseudo-label
@@
8:  $c_{i}\leftarrow\frac{1}{d_{i}+\epsilon}$ $\triangleright$ Compute confidence score
9:  $\mathit{confidences}.\text{append}((x_{i},\hat{y}_{i},c_{i}))$ $\triangleright$ Store sample, pseudo-label, and confidence
@@
11:Sort $\mathit{confidences}$ by $c_{i}$ in descending order
@@
IV-D2 Training Network with Local Alignment
In Step 2, we train our network using the loss function consisting of three components, presented as:
@@ IV-D2 Training Network with Local Alignment
where $\alpha_{1},\alpha_{2}$ represent the weights. In the loss function, firstly, $L_{cls}$ represents the traditional cross-entropy loss function, which utilizes data solely from the training set. It is used to train the model to learn the basic classification capacity. Secondly, $L_{MMD}^{local}$ represents the local MMD, which calculates the MK-MMD between the help set and the training set within each category:
@@
where $M$ represents the total category amounts, $E$ is the embedding result of ResNet, $K$ is the given kernel list, and $Y$ is the sample label. It’s used to realize our local alignment. Additionally, we add an extra global MMD, which computes the MK-MMD between the entire help set and training set, following the conventional MMD approach, represented as:
@@
The early stop strategy is outlined in Algorithm 2. During the training, we keep track of the highest accuracy and the lowest value of the loss function in the support set. Each time the accuracy or the loss value is improved, we save the corresponding model parameters. When the epoch reaches the minimum training epoch $e^{min}$, if the validation accuracy does not improve and the validation loss does not decrease for consecutive $e^{threshold}$ epochs, the model is regarded as converged. Specifically, in our method, when we reach the epoch $e^{min}$, we incorporate two different factors to decrease the historical highest accuracy and increase the historical largest loss. This is because in our experiments, the model sometimes achieves an artificially high accuracy and low loss at the beginning of training, which is not a desirable outcome. The underlying reason for this is the small size of the validation set. Especially in the one-shot scenario, there is only a single sample for each category in the validation set (i.e., the support set), which can lead to relatively high fluctuations in loss and accuracy during training. Therefore, by utilizing these factors, we aim to increase the likelihood of the neural network preserving the best parameters after reaching $e^{min}$.
@@
$X^{train},Y^{train}]$: Training data and labels
$X^{support},Y^{support}]$: Support set for validation
$X^{help},Y^{help}]$: Help set for domain adaptation
@@
[…]
@@
8:  $E\leftarrow\text{ResNet}(X^{train})$ $\triangleright$ Extract features
9:  $\hat{Y}\leftarrow\text{MLP}(E)$ $\triangleright$ Compute predictions
@@
11:  $l^{train}\leftarrow\mathcal{L}(\hat{Y},Y^{train})$ $\triangleright$ Eq. 11
12:  $\theta\leftarrow\theta-\eta\nabla_{\theta}l^{train}$ $\triangleright$ Parameter update
13:  Validation on Support Set:
14:  $l\leftarrow\text{CrossEntropy}(X^{support},Y^{support})$ $\triangleright$ Validation loss
15:  $a\leftarrow\text{Accuracy}(X^{support},Y^{support})$ $\triangleright$ Validation accuracy
@@
20:   $\theta^{best}\leftarrow\theta$ $\triangleright$ Save improved parameters
@@
28:   $\theta^{best}\leftarrow\theta$ $\triangleright$ Save improved parameters
@@
34:   $loss^{best}\leftarrow loss^{best}\cdot\alpha$ $\triangleright$ Relax loss threshold
35:   $acc^{best}\leftarrow acc^{best}\cdot\beta$ $\triangleright$ Relax accuracy threshold
36:   $e^{acc},e^{loss}\leftarrow 1$ $\triangleright$ Reset counters
@@
43:Return $\theta^{best}$ $\triangleright$ Return best parameters found
V Experiment
V-A Dataset Description
In our experiment, we utilize two datasets: the publicly available WiGesture dataset [4] and a novel dataset named WiFall, which is proposed along with this paper. Both datasets are collected using the ESP32-S3 device, which operates in the 2.4 GHz Wi-Fi band-the default central frequency supported by the hardware and share the same data format, with a sample rate of 100Hz, 1 antenna, and 52 subcarriers. For our experiment, we divide the data into 1-second samples, resulting in each sample having a length of 100 data points.
V-A1 WiGesture Dataset
@@ V-B Experiment Setup
Selected Proportion in Help Set : $p\%$ 50%
@@
$e^{min},e^{max},e^{threshold},\alpha,\beta]$
@@
To evaluate the efficiency of our method, we compare it with several representative cross-domain methods in the fields of Wi-Fi sensing and machine learning. For fairness, we have made every effort to replicate other works according to the descriptions provided in their respective papers. For any missing settings in these papers, we adopt the same settings as our method. In the field of Wi-Fi sensing, most works are based on the feature learning (learning-based) paradigm. We choose the four works mentioned before for comparison: Ding et al. [18], AutoFi [19], Yang et al. [20], and CrossFi [22], which represent different powerful feature extractors and similarity calculation techniques. Even some works combined with DAL methods, such as Yang et al. [20] and CrossFi [22], which utilize the MK-MMD loss, do not rely solely on it. To further illustrate the performance of the DAL methods, we have chosen several popular methods in machine learning, including adversarial learning-based DANN [31] and ADDA [46], as well as feature-space alignment methods such as MK-MMD [43] and GFK [29]. Since these methods do not depend on specific network structures, we use the same ResNet architecture as in this paper. Additionally, we employ the same kernel functions in MK-MMD as in our KNN-MMD method. Furthermore, we also compare the methods of Tian et al. [39] and EEG [40], which also discuss local alignment methods, with detailed descriptions provided in Section III. All these DAL methods fall under the zero-shot category, meaning they only require a labeled training set from the source domain and an unlabeled testing set, without the necessity for a labeled support set.
V-C Experiment Result
@@
V-D Comparison with Traditional DA Methods
@@
V-D1 Metric-based Method
For metric-based methods like KNN [25], the biggest problem is that they are heavily influenced by the quality of the support set. As shown in Table VI, we reduced the data dimension in the target domain to $d$ using UMAP and tested the KNN performance in an $n$-shot scenario with $k$ nearest neighbors. Each scenario was tested 3 times, and the table shows the accuracy range. It can be seen that the accuracy fluctuates greatly, which may pose challenges for the practical application of this method. In Fig. 9, it can be seen that our KNN-MMD method does not exhibit such large fluctuations. Moreover, in Table VI, it’s noticeable that our KNN-MMD is not very sensitive to the hyper-parameters.
@@ V-D1 Metric-based Method
V-D2 Learning-based Method
@@
V-D3 DAL Method
As mentioned before, the global alignment operation of traditional DAL methods cannot guarantee the model performance. In Fig. 11, we illustrate the UMAP result of different methods, including: ResNet (no alignment), MK-MMD (global alignment), KNN-MMD (shot number=1), and Theoretical Upper Bound (local alignment).
It can be seen that the performance of the global alignment method MK-MMD does not differ much from ResNet, which even has no alignment operation. Even though they can split samples in the source domain (training set) well, achieving an accuracy over 99%, they cluster samples in the target domain (testing set) chaotically. In contrast, our KNN-MMD still performs well in the target domain. As shown in Fig. 11c, it successfully aligns samples from the source domain and target domain within each category. Furthermore, in Fig. 11d, we show the upper bound performance of KNN-MMD, where we provide the ground truth label of the help set instead of pseudo-labels. We notice that the cluster performance and accuracy of KNN-MMD and the upper bound do not have a significant difference, which proves the efficiency of our help set construction method. Finally, even though our early stop method can be used in most DAL methods, as the support set does not directly participate in training, limited by the model principle, even if we make it stop at the point with the best testing accuracy, it still has a poor performance, as shown in Fig. 10b.
@@ V-E Ablation Study
V-F Expanded Experiment
V-F1 Evaluation on WiDar3.0
@@ V-F Expanded Experiment
To further illustrate the efficiency of our KNN-MMD, we conduct cross-domain experiments using the WiDar3.0 dataset [48], which includes both cross-people and cross-room gesture recognition scenarios. In WiDar3.0, the Body-coordinate Velocity Profile (BVP) extracted from CSI is provided, with the dimension of $t\times 20\times 20$, where $t$ represents the time length. To standardize the data dimensions, we utilize only the first 28 frames of each sample and filter out those samples that are shorter than 28 frames. For our experiment, we select the subsets as follows: (i) Cross-People Scenario: For the cross-person gesture task, we utilize data collected on November 9, 2018, which includes six gestures: Push&Pull, sweep, clap, slide, draw, zigzag, and draw-N. We designate User 1 as the source domain and User 2 as the target domain. Both the source and target domains contain 3,000 samples. (ii) Cross-Room Scenario: For the cross-room gesture task, we utilize data collected on November 21 and November 27, 2018, from User 2, which includes six gestures: slide, draw-O, draw zigzag, draw-N, draw-triangle, and draw-rectangle. We designate Room 1 as the source domain and Room 2 as the target domain. The source domain contains 3,000 samples, while the target domain has 750 samples.
The experimental results are shown in Fig. 12, where our KNN-MMD demonstrates excellent performance in both cross-people and cross-room tasks, slightly outperforming CrossFi and significantly outperforming other benchmarks. Since AutoFi [19] only supports CSI input due to its pre-training method, we have omitted it from this section.
(a) Cross-People Scenario
(b) Cross-Room Scenario
Figure 12: Experiment Results on WiDar3.0 [48]
V-F2 Influence of Model Input
In many Wi-Fi sensing studies, processed CSI signals, such as Time of Flight (ToF), Angle of Arrival (AoA), Attenuation, and Doppler Frequency Shift (DFS), are commonly used instead of the original CSI. These processed signals serve as important feature extraction methods in model-based approaches. However, this practice remains controversial in the field of deep learning. While data pre-processing methods can help extract meaningful features, they may also lead to the loss of information present in the original data. As noted in [48], DFS contains the most information about velocity distribution, offering advantages over other features. Additionally, Micro-Doppler has been successfully used in OFDM radar sensing tasks and has also shown promising results in Wi-Fi sensing [49]. To investigate this, we compare the performance of KNN-MMD on the WiGesture dataset using both original CSI signals, DFS, and Micro-Doppler. Specifically, DFS is obtained by estimating the frequency shift from the phase differences of consecutive CSI samples, which primarily captures macro motions and change. For a CSI stream denoted as $H(t)=|H(t)|e^{j\angle H(t)}$, the phase difference between consecutive samples is used to estimate the Doppler shift $f_{d}$ via: $f_{d}=\frac{1}{2\pi}\frac{\mathrm{d}(\angle H(t))}{\mathrm{d}t}$. Micro-Doppler, in contrast, is extracted by applying time-frequency analysis such as the Short-Time Fourier Transform (STFT) to the CSI series, which reveals finer-grained motion patterns. Shown as Fig. 13, the results indicate that, compared to original CSI input, DFS can improve performance in people identification but may lead to lower performance in gesture recognition. Conversely, Micro-Doppler demonstrates an opposite trend. The underlying reasons for these observations may be complex; while data pre-processing can provide prior knowledge to the network, it may also result in potential information loss from the original data. Nonetheless, the overall performance of our method does not vary significantly among different inputs, illustrating its robustness. In practice, the choice of input can be tailored to the specific task and available computational resources.